An investigator compares the durability of two different compounds used in the manufacture of a certain automobile brake lining. A sample of 138 brakes using Compound 1 yields an average brake life of 30,636 miles. A sample of 111 brakes using Compound 2 yields an average brake life of 43,730 miles. Assume that the population standard deviation for Compound 1 is 3899 miles, while the population standard deviation for Compound 2 is 1161 miles. Determine the 95% confidence interval for the true difference between average lifetimes for brakes using Compound 1 and brakes using Compound 2.

Step 1 of 3: Find the point estimate for the true difference between the population means.

Step 2 of 3: Calculate the margin of error of a confidence interval for the difference between the two population means. Round your answer to six decimal places.

Step 3 of 3: Construct the 95% confidence interval. Round your answers to the nearest whole number.
Answer #1

1)

Point estimate = 1 - 2 = 30636 - 43730 = -13094

2)

Margin of error E = Z * sqrt ( 1 / n1 + 2 / n2 )

= 1.96 * sqrt ( 3899 / 138 + 1161 / 111)

= 12.195088

3)

95% confidence interval is

(1 - 2 ) - E < 1 - 2 < (1 - 2 ) + E

-13094 - 12.195088 < 1 - 2 < -13094 + 12.195088

-13106 < 1 - 2 < -13082

95% CI is ( -13106 , -13082 )